# A 13/6-Approximation for Strip Packing via the Bottom-Left Algorithm

### v1

@@ 1 Introduction
The Strip Packing problem is a generalization of both Bin Packing and Load Balancing ($P||C_{\text{max}}$), which implies that the problem is NP-hard [14] and that there cannot exist a polynomial time algorithm with absolute approximation ratio $3/2-\varepsilon$, unless P=NP. Additionally, the problem is known to be strongly NP-hard [8].
These hardness results provide strong motivation for the development of efficient approximation algorithms. The earliest of such algorithms for Strip Packing is the Bottom-Left Algorithm (called BL algorithm for short), introduced in $1980$ by Baker, Coffman, and Rivest [1]. This algorithm is conceptually simple yet powerful: it processes the input rectangles in a predetermined order, placing each at the lowest possible position in the strip and breaking ties by choosing the leftmost position. Despite its simplicity, the BL algorithm laid the foundation for decades of research into geometric packing heuristics and many open questions surrounding the algorithm remain. In their pioneering work, Baker et al.[1] proved that the BL algorithm achieves an absolute approximation ratio of $3$, and no better bound for this algorithm has been established since. Furthermore, Chazelle [3] established a quadratic runtime implementation of the BL algorithm.
@@
As the performance of the BL algorithm strongly depends on the predetermined order of the rectangles, it is tempting to think that among the $n!$ possible orderings of $n$ rectangles, at least one of them produces an optimal packing. Unfortunately, this is not the case: even under the best possible ordering, the absolute approximation ratio cannot surpass $4/3-\varepsilon$ [11], a fact already foreshadowed in earlier works [1, 2]. Nevertheless, the BL algorithm is a $3$-approximation when rectangles are ordered by decreasing width (using an arbitrary order for rectangles with the same width) [1].
@@
Table 1: Upper bounds on the approximation ratio of the Bottom-Left Algorithm for different orderings of squares and rectangles. Results with a star have a matching lower bound.
@@
Related Work. Kenyon and Rémila [15] proved that there exists an asymptotic fully polynomial time approximation scheme for Strip Packing. Their result has been sharpened in [13, 20]. Beyond classical Strip Packing, the Bottom-Left Algorithm has also been studied in other contexts. In Multiple Strip Packing, where there are several strips of different widths, Zhuk [21] constructs a $10$-approximation that combines an online algorithm for assigning the rectangles to strips, with the BL algorithm using decreasing-width order to place the rectangles within the strip. Furthermore, Fekete, Kamphans, and Schweer [6] considered the online variant of Square Strip Packing with Tetris and gravity constraints, proving a competitive ratio of $3.5$. The methods developed in this paper may also be useful in these contexts, offering the potential for improved bounds or deeper understanding of the BL algorithm in its various extensions.
@@ 2.2 The horizontal strip partition
To be more precise, let $\mathcal{R}=\{r_{1},\dots,r_{n}\}$ and assume that the BL algorithm places $r_{i}$ before $r_{j}$ if $i<j$. We define the horizontal strip partition as the partition of the strip $0,W]\times[0,h_{\text{BL}}(\mathcal{R},W)]$ into the regions
@@ 2.3 Covering proper horizontal lines
For this, we want to analyze the fraction of a horizontal line in the strip that is occupied by rectangles. We call a horizontal line proper if it does not intersect the top or bottom face of rectangles. We do not need to consider lines that are not proper, because their total area is zero. A proper horizontal line alternates between parts that are occupied by rectangles and parts that are unoccupied. An unoccupied gap is a maximal connected unoccupied part on the line.
@@ 3 The 13/6-approximation
This section is devoted to proving that the Bottom-Left Algorithm is a $13/6$-approximation when placing the rectangles in the so-called $\mathcal{FQW}$-ordering. We start in Section 3.1 by constructing this novel ordering that is based on the $\mathcal{FQW}$-partition of the rectangles, and establish several key properties of the partition. After that, in Section 3.2, we use the horizontal strip partition to derive lower bounds on the area occupied by rectangles in the BL packing. Most regions in this partition are at least half occupied by rectangles, though a few require special attention, being in total occupied only up to a $5/12$ fraction. Section 3.3 develops tools to distinguish different types of horizontal lines, which leads to the construction of a quadratic program to bound the maximum occupied area in the special regions. This analysis culminates in the proof of Theorem 1 in Section 3.4, followed by final remarks on our approach in Section 3.5.
@@ 3.2 Setup of analysis
We index the rectangles from the instance $(\mathcal{R},W)$ by $r_{1},\dots,r_{n}$ such that the BL algorithm places $r_{i}$ before $r_{j}$ if $i<j$. For $a=|\mathcal{F}|$ and $b=|\mathcal{Q}|$, it now holds that $\mathcal{F}=\{r_{1},\dots,r_{a}\}$$\mathcal{Q}=\{r_{a+1},\dots,r_{a+b}\}$ and $\mathcal{W}=\{r_{a+b+1},\dots,r_{n}\}$. See Figure 1.
@@
It remains to study the regions $H_{i}$ for $i\in\{a+2,\dots,a+b\}$. For this, we define the left rectangle of $\mathcal{Q}$ to be the first rectangle $r_{L}$ from $\mathcal{Q}$ that touches the left strip boundary (in the example shown in Figure 1 the rectangle $r_{8}$ is the rectangle $r_{L}$). If no rectangle in $\mathcal{Q}$ touches the left strip boundary, then we set $r_{L}:=r_{T}$. In Section 3.3 we will develop a technique to prove that the union $H_{a+2}\cup\cdots\cup H_{L}$ is at least for a $5/12$-fraction occupied by rectangles. But first, we consider the regions $H_{L+1},\dots,H_{a+b}$. Since the rectangles in $\mathcal{Q}$ are ordered by decreasing width we can apply Lemma 3.
@@ 3.3 The space $H_{a+2}\cup\cdots\cup H_{L}$
A proper horizontal line in the regions $H_{a+2},\dots,H_{L}$ might be less than half occupied prior to placing the first rectangle above the line, because, contrary to before, the leftmost part of the line is either occupied by rectangles from $\mathcal{F}$ or is unoccupied. However, if the line intersect $k$ rectangles from $\mathcal{Q}$ prior to placing the first rectangle above the line, then, according to Lemma 3, the line is at least a $k/(2k+1)\geq 1/3$-fraction occupied. The central idea of this section is to establish a bound on the number of lines intersecting a fixed number of rectangles, which will subsequently imply that the union of regions $H_{a+2}\cup\cdots\cup H_{L}$ is occupied by at least a $5/12$ fraction.
@@
Let $\mathcal{L}^{k}$ be the set of all proper horizontal lines of order $k$ inside of the space $H_{a+2}\cup\cdots\cup H_{L}$. Define $\mathcal{R}^{k}=\{r\in\mathcal{Q}\mid\exists\ell\in\mathcal{L}^{k}:r={LM}(\ell)\}$, that is, the set of all rectangles in $\mathcal{Q}$ that is the leftmost rectangle of some line of order $k$. Furthermore, we denote by ${LM}^{k}$ the leftmost rectangle from $\mathcal{R}^{k}$, and if there are multiple such rectangles, then let ${LM}^{k}$ be the highest among them.
@@
Figure 2: Example showing that the lower horizontal line $\ell$ of order 22 is at least half occupied. Blue rectangles are in $\mathcal{F}$ and brown ones are in $\mathcal{Q}$.
@@
The space $H_{a+2}\cup\cdots\cup H_{L}$ is at least a $5/12$ fraction occupied by rectangles.
@@
By Corollary 15 the space $H_{a+2}\cup\cdots\cup H_{L}$ is maximally unoccupied when $|\mathcal{R}^{k}|\leq 1$ for all $k$, because only lines of order $k$ intersecting the leftmost rectangle of $\mathcal{R}^{k}$ can be $(k+1)/(2k+1)$ unoccupied, while other lines of order $k$ are only at most half unoccupied.
Define $\alpha_{k}=\sup\{d(\ell,\ell^{\prime})\mid\ell,\ell^{\prime}\in\mathcal{L}^{k}\text{ s.t. }{LM}^{k}\in T(\ell)\cap T(\ell^{\prime})\}$, that is, the largest distance between two lines of order $k$ that both intersect the leftmost rectangle of $\mathcal{R}^{k}$. From Lemma 14 it follows that the only lines of order $k$ that can be less than half occupied, are the ones in this supremum. It holds that $\alpha_{k}\leq h_{{LM}^{k}}\leq\frac{1}{2}h_{\text{OPT}}$ by Lemma 4. Furthermore, by Lemma 11, the height of the space $H_{a+2}\cup\cdots\cup H_{L}$ is at most $y_{r_{L}}-y_{r_{a+1}}\leq h_{\text{max}}\leq h_{\text{OPT}}$, because $y_{r_{a+1}}=h_{r_{B}}$, where $r_{B}$ is the leftmost bottom supporter of $r_{a+1}$. Hence we also have that $\sum_{k=1}^{\infty}\alpha_{k}\leq h_{\text{OPT}}$, and to maximize the unoccupied space, we can assume that in the worst case this is an equality. If we normalize such that $h_{\text{OPT}}=1$, then we have the linear constraints $\alpha_{k}\leq\frac{1}{2}$ and $\sum_{k=1}^{\infty}\alpha_{k}=1$.
Next, consider $\ell\in\mathcal{L}^{k}$ that is less than half occupied. As Figure 3 suggests, this line might intersect some rectangles $r_{1},\dots,r_{z}$ from $\mathcal{F}$ and it intersects $k$ rectangles $r_{i_{1}},\dots,r_{i_{k}}$ from $\mathcal{Q}$ prior to placing the first rectangle above $\ell$. Define $x_{\mathcal{F}}^{\ell}=x_{r_{z}}+w_{r_{z}}$ if $r_{z}$ exists, and otherwise define $x_{\mathcal{F}}^{\ell}=0$. By Lemma 2 at this moment, each gap between rectangles has width strictly less than the width of the smallest rectangle from $\mathcal{Q}$ that intersects the line $\ell$. Hence, there exists an $x$-coordinate $x_{\ell}\in[x_{\mathcal{F}}^{\ell},x_{r_{i_{1}}}]$ such that the line is exactly half occupied on the right of $x_{\ell}$. Define the length of the remaining gap on the line by $\beta^{\ell}=x_{\ell}-x_{\mathcal{F}}^{\ell}$.
@@
Figure 3: Example of a horizontal line $\ell$ of order $2$. The right of $x_{\ell}$ is exactly half occupied, the remaining gap on the left of $x_{\ell}$ is defined as $\beta^{\ell}$.
Furthermore, for all lines of order $k$ that are less than half occupied, consider the largest such $\beta^{\ell}$-gap
$\beta_{k}=\max\{\beta^{\ell}\mid\ell\in\mathcal{L}^{k}\text{ is less than $1/2$ occupied}\}.$
Normalize such that $W=1$. It holds that the gap $\beta_{k}$ is upper bounded by $\frac{1}{2k+1}$, which is tight if $r_{z}$ does not exist and $x_{\mathcal{F}}^{\ell}=x_{r_{i_{1}}}$. Moreover, it holds that $\sum_{k=1}^{\infty}\beta_{k}\leq 1$, because for different $k$ the downward projections of the gaps defined by $\beta_{k}$ are disjoint by Lemma 16.
Next we describe the relation between the amount of unoccupied space and the values $\alpha$ and $\beta$. A line $\ell$ of order $k$ has exactly $\beta^{\ell}+\frac{1}{2}\left(1-x_{\ell}\right)$ unoccupied space (see Figure 3). Clearly it holds that $\beta^{\ell}\leq\beta_{k}$. Furthermore, we have $x_{\ell}\geq\sum_{j=k}^{\infty}\beta_{j}$, hence $\ell$ has at most $\beta_{k}+\frac{1}{2}\left(1-\sum_{j=k}^{\infty}\beta_{j}\right)$ unoccupied space. At most $\alpha_{k}$ of the lines of order $k$ are more than half unoccupied by Lemma 14. Hence the maximum amount of unoccupied space in $H_{a+2}\cup\cdots\cup H_{L}$ is a solution to the following quadratic infinite program
$\max\quad\quad$ $\sum_{k=1}^{\infty}\left(\alpha_{k}\beta_{k}+\frac{1}{2}\alpha_{k}\left(1-\sum_{j=k}^{\infty}\beta_{j}\right)\right)\quad\text{such that},$
$0\leq\alpha_{k}\leq\frac{1}{2}\quad\text{for all }k\in\mathbb{N},$
$\sum_{k=1}^{\infty}\alpha_{k}=1,$
$0\leq\beta_{k}\leq\frac{1}{2k+1}\quad\text{for all }k\in\mathbb{N},$
$\sum_{k=1}^{\infty}\beta_{k}\leq 1.$
@@
$\max\quad\quad\frac{1}{2}$ $\sum_{k=1}^{\infty}\alpha_{k}\gamma_{k}\quad\text{where }\gamma_{k}=\beta_{k}+1-\sum_{j=k+1}^{\infty}\beta_{j}.$
Let $k_{1}$ and $k_{2}$ be different indices such that $\gamma_{k_{1}}\geq\gamma_{k_{2}}\geq\gamma_{k}$ for all $k\in\mathbb{N}\setminus\{k_{1}\}$. Then setting $\alpha_{k_{1}}=\alpha_{k_{2}}=\frac{1}{2}$ and $\alpha_{k}=0$ for $k\in\mathbb{N}\setminus\{k_{1},k_{2}\}$ maximizes the objective (provided that $\gamma_{k_{1}},\gamma_{k_{2}}\geq 0$). Hence it remains to maximize the linear objective $\frac{1}{4}(\gamma_{k_{1}}+\gamma_{k_{2}})$. Observe that for $i\leq j$ it holds that
$\gamma_{i}+\gamma_{j}=\beta_{i}+\beta_{j}+2-\sum_{k=i+1}^{\infty}\beta_{k}-\sum_{k=j+1}^{\infty}\beta_{k}=\beta_{i}+2-\sum_{k=i+1}^{j-1}\beta_{k}-2\sum_{k=j+1}^{\infty}\beta_{k}.$
This is maximal if $\beta_{k}=0$ for all $k\in\{i+1,\dots,j-1,j+1,\dots\}$. Then it holds that $\gamma_{i}+\gamma_{j}=\beta_{i}+2$, which is maximal if $i=1$, and $\beta_{1}=\frac{1}{3}$. From this it follows that the value of an optimum solution of the infinite quadratic program is $\frac{1}{4}(\frac{1}{3}+2)=\frac{7}{12}$. Thus the space $H_{a+2}\cup\cdots\cup H_{L}$ is at least $5/12$ occupied. Moreover, notice that this optimal value is obtained for $\alpha_{1}=\alpha_{2}=\frac{1}{2}$$\alpha_{k}=0$ for $k\geq 3$$\beta_{1}=\frac{1}{3}$ and $\beta_{k}=0$ for all $k\geq 2$. ∎

### v2

@@ 1 Introduction
The Strip Packing problem is a generalization of both Bin Packing and Load Balancing ($P||C_{\text{max}}$), which implies that the problem is NP-hard [14] and that there cannot exist a polynomial-time algorithm with absolute approximation ratio $3/2-\varepsilon$, unless P=NP. Additionally, the problem is known to be strongly NP-hard [8].
These hardness results provide strong motivation for the development of efficient approximation algorithms. The earliest of such algorithms for Strip Packing is the Bottom-Left Algorithm (called the BL algorithm for short), introduced in $1980$ by Baker, Coffman, and Rivest [1]. This algorithm is conceptually simple yet powerful: it processes the input rectangles in a predetermined order, placing each at the lowest possible position in the strip and breaking ties by choosing the leftmost position. Despite its simplicity, the BL algorithm laid the foundation for decades of research into geometric packing heuristics and many open questions surrounding the algorithm remain. In their pioneering work, Baker et al.[1] proved that the BL algorithm achieves an absolute approximation ratio of $3$, and no better bound for this algorithm has been established since. Furthermore, Chazelle [3] established a quadratic runtime implementation of the BL algorithm.
@@
As the performance of the BL algorithm strongly depends on the predetermined order of the rectangles, it is tempting to think that among the $n!$ possible orderings of $n$ rectangles, at least one of them produces an optimal packing. Unfortunately, this is not the case: even under the best possible ordering, the absolute approximation ratio of the BL algorithm cannot surpass $4/3-\varepsilon$ [11], a fact already foreshadowed in earlier works [1, 2]. Nevertheless, the BL algorithm is a $3$-approximation when rectangles are ordered by decreasing width (using an arbitrary order for rectangles with the same width) [1].
@@
Decreasing area $2^{\star}$ (as above) unbounded (Appendix)
@@
Increasing height 33 (as above) $\geq 3$ [22]
Increasing area 33 (as above) unbounded (Appendix)
@@
Table 1: Upper/lower bounds on the approximation ratio of the Bottom-Left Algorithm for different orderings of squares and rectangles. Results with a star have a matching lower bound.
@@
Related Work.
Kenyon and Rémila [15] proved that there exists an asymptotic fully polynomial-time approximation scheme for Strip Packing. Their result has been sharpened in [13, 20]. Beyond classical Strip Packing, the Bottom-Left Algorithm has also been studied in other contexts. In Multiple Strip Packing, where there are several strips of different widths, Zhuk [21] constructs a $10$-approximation that combines an online algorithm for assigning the rectangles to strips, with the BL algorithm using decreasing-width order to place the rectangles within the strip. Furthermore, Fekete, Kamphans, and Schweer [6] considered the online variant of Square Strip Packing with Tetris and gravity constraints, proving a competitive ratio of $3.5$. The methods developed in this paper may also be useful in these contexts, offering the potential for improved bounds or deeper understanding of the BL algorithm in its various extensions.
@@ 2.2 The horizontal strip partition
Formally, let $\mathcal{R}=\{r_{1},\dots,r_{n}\}$ such that the BL algorithm places $r_{i}$ before $r_{j}$ if $i<j$. We define the horizontal strip partition of a BL packing as the partition of the strip $0,W]\times[0,h_{\text{BL}}(\mathcal{R},W)]$ into the regions
@@ 2.3 Covering proper horizontal lines
For this, we want to analyze the fraction of a horizontal line in the strip that is occupied by rectangles. We call a horizontal line proper if it does not intersect the top or bottom face of any rectangle. We do not need to consider lines that are not proper, because their total area is zero. A proper horizontal line alternates between parts that are occupied by rectangles and parts that are unoccupied. An unoccupied gap is a maximal connected unoccupied part on the line.
@@ 3 The 13/6-approximation
This section is devoted to proving that the Bottom-Left Algorithm is a $13/6$-approximation when placing the rectangles in the so-called $\mathcal{FQW}$-ordering. We start in Section 3.1 by constructing this novel ordering that is based on the $\mathcal{FQW}$-partition of the rectangles, and establish several key properties of the partition. After that, in Section 3.2, we use the horizontal strip partition to derive lower bounds on the area occupied by rectangles in the BL packing. Most regions in this partition are at least half occupied by rectangles, though a few require special attention, being in total a little less than half occupied. Section 3.3 develops tools to distinguish different types of horizontal lines, which leads to the construction of a quadratic program to bound the maximum occupied area in the special regions. This analysis culminates in the proof of Theorem 1 in Section 3.4, followed by final remarks on our approach in Section 3.5.
@@ 3.2 Setup of analysis
We index the rectangles from the instance $(\mathcal{R},W)$ by $r_{1},\dots,r_{n}$ such that the BL algorithm places $r_{i}$ before $r_{j}$ if $i<j$. Furthermore, we denote $a=|\mathcal{F}|$ and $b=|\mathcal{Q}|$, which implies that $\mathcal{F}=\{r_{1},\dots,r_{a}\}$$\mathcal{Q}=\{r_{a+1},\dots,r_{a+b}\}$ and $\mathcal{W}=\{r_{a+b+1},\dots,r_{n}\}$ (see Figure 1).
@@
It remains to study the regions $H_{i}$ for $i\in\{a+2,\dots,a+b\}$. For this, we define the left rectangle of $\mathcal{Q}$ to be the first rectangle $r_{L}$ from $\mathcal{Q}$ that touches the left strip boundary (in the example shown in Figure 1 the rectangle $r_{8}$ is the rectangle $r_{L}$). If no rectangle in $\mathcal{Q}$ touches the left strip boundary, then we set $r_{L}:=r_{T}$. In Section 3.3 we will develop a technique to prove that the union $H_{a+2}\cup\cdots\cup H_{L}$ is almost half occupied by rectangles. But first, we consider the regions $H_{L+1},\dots,H_{a+b}$. Since the rectangles in $\mathcal{Q}$ are ordered by decreasing width we can apply Lemma 3.
@@ 3.3 The space $H_{a+2}\cup\cdots\cup H_{L}$
A proper horizontal line in the regions $H_{a+2},\dots,H_{L}$ might be less than half occupied prior to placing the first rectangle above the line, because, contrary to before, the leftmost part of the line is either occupied by rectangles from $\mathcal{F}$ or is unoccupied. However, if the line intersects $k$ rectangles from $\mathcal{Q}$ prior to placing the first rectangle above the line, then, according to Lemma 3, the line is at least a $k/(2k+1)\geq 1/3$-fraction occupied. The central idea of this section is to establish a bound on the number of lines intersecting a fixed number of rectangles, which will subsequently imply that the union of the regions $H_{a+2}\cup\cdots\cup H_{L}$ is occupied by at least the amount $\frac{1}{2}hW-\frac{1}{12}h_{\text{OPT}}W$ where $h=y_{r_{L}}-y_{r_{a+1}}$ is the height of the union of these regions.
@@
Let $\mathcal{L}^{k}$ be the set of all proper horizontal lines of order $k$ inside of the space $H_{a+2}\cup\cdots\cup H_{L}$. Define $\mathcal{R}^{k}=\{r\in\mathcal{Q}\mid\exists\ell\in\mathcal{L}^{k}:r={LM}(\ell)\}$, that is, the set of all rectangles in $\mathcal{Q}$ that are the leftmost rectangle of some line of order $k$. Furthermore, we denote by ${LM}^{k}$ the leftmost rectangle from $\mathcal{R}^{k}$, and if there are multiple such rectangles, then let ${LM}^{k}$ be the highest among them.
@@
Figure 2: Example showing that the lower horizontal line $\ell$ of order $2$ is at least half occupied. Blue rectangles are in $\mathcal{F}$ and brown ones are in $\mathcal{Q}$.
@@
The union of the regions $H_{a+2}\cup\cdots\cup H_{L}$ consists of at least $\frac{1}{2}hW-\frac{1}{12}h_{\text{OPT}}W$ occupied space with $h=y_{r_{L}}-y_{r_{a+1}}$.
@@
By Corollary 15 the space $H_{a+2}\cup\cdots\cup H_{L}$ is maximally unoccupied when we have for all $k$ that $|\mathcal{R}^{k}|\leq 1$, because only lines of order $k$ intersecting the leftmost rectangle of $\mathcal{R}^{k}$ can be $(k+1)/(2k+1)$ unoccupied, while other lines of order $k$ are only at most half unoccupied.
Define $\alpha_{k}=\sup\{d(\ell,\ell^{\prime})\mid\ell,\ell^{\prime}\in\mathcal{L}^{k}\text{ s.t. }{LM}^{k}\in T(\ell)\cap T(\ell^{\prime})\}$, that is, the largest distance between two lines of order $k$ that both intersect the leftmost rectangle of $\mathcal{R}^{k}$. If no line of order $k$ exists, then define $\alpha_{k}=0$. From Lemma 14 it follows that all lines of order $k$ that are less than half occupied are considered in this supremum. It holds that $\alpha_{k}\leq h_{{LM}^{k}}\leq\frac{1}{2}h_{\text{OPT}}$ by Lemma 4. Furthermore, the height of the space $H_{a+2}\cup\cdots\cup H_{L}$ is $h$ by definition, thus $\sum_{k=1}^{\infty}\alpha_{k}=h$. Notice that for an instance consisting of $n$ rectangles, it holds that $\alpha_{k}=0$ for any $k>n$, hence $\sum_{k=1}^{\infty}\alpha_{k}$ equals the finite sum $\sum_{k=1}^{n}\alpha_{k}$. In conclusion, we have the linear constraints $\alpha_{k}\leq\frac{1}{2}h_{\text{OPT}}$ for every $k$ and $\sum_{k=1}^{\infty}\alpha_{k}=h$.
Next, consider $\ell\in\mathcal{L}^{k}$ that is less than half occupied. This line might intersect some rectangles $r_{1},\dots,r_{z}$ from $\mathcal{F}$ and it intersects $k$ rectangles $r_{i_{1}},\dots,r_{i_{k}}$ from $\mathcal{Q}$ prior to placing the first rectangle above $\ell$ (see Figure 3). We define $x_{\mathcal{F}}^{\ell}=x_{r_{z}}+w_{r_{z}}$ if $r_{z}$ exists, and otherwise we define $x_{\mathcal{F}}^{\ell}=0$. By Lemma 2 at this moment, each gap between rectangles has width strictly less than the width of the smallest rectangle from $\mathcal{Q}$ that intersects the line $\ell$. Hence, there exists an $x$-coordinate $x_{\ell}\in[x_{\mathcal{F}}^{\ell},x_{r_{i_{1}}}]$ such that the line $\ell$ is exactly half occupied to the right of $x_{\ell}$. Define the length of the remaining gap on the line by $\beta^{\ell}=x_{\ell}-x_{\mathcal{F}}^{\ell}$.
@@
Figure 3: Example of a horizontal line $\ell$ of order $2$. To the right of $x_{\ell}$ the line $\ell$ is exactly half occupied prior to placing the first rectangle above $\ell$, the remaining gap between $x_{\mathcal{F}}^{\ell}$ and $x_{\ell}$ is defined as $\beta^{\ell}$.
Furthermore, for all lines of order $k$ that are less than $W/2$ occupied, consider the largest such $\beta^{\ell}$-gap
$\beta_{k}=\max\{\beta^{\ell}\mid\ell\in\mathcal{L}^{k}\text{ is less than $W/2$ occupied}\}.$
Again define $\beta_{k}=0$ if $\mathcal{L}^{k}$ is empty. It holds that the gap $\beta_{k}$ is upper bounded by $\frac{1}{2k+1}W$, because by Lemma 3 a line of order $k$ is at most $\frac{k}{2k+1}W$ occupied and each rectangle on the line is larger than any unoccupied gap on the line. Moreover, it holds that $\sum_{k=1}^{\infty}\beta_{k}\leq W$, because for different values of $k$ the downward projections of the gaps defined by $\beta_{k}$ are disjoint by Lemma 16.
Next we describe the relation between the amount of unoccupied space and the values $\alpha$ and $\beta$. A line $\ell$ of order $k$ has exactly $\beta^{\ell}+\frac{1}{2}\left(W-x_{\ell}\right)$ unoccupied space (see Figure 3). By definition it holds that $\beta^{\ell}\leq\beta_{k}$. Furthermore, we have $x_{\ell}\geq\sum_{j=k}^{\infty}\beta_{j}$, because the rightmost point of the $\beta_{j+1}$-gap is left of the leftmost point of the $\beta_{j}$-gap for all $j$ by Lemma 16. Therefore, $\ell$ has at most $\beta_{k}+\frac{1}{2}\left(W-\sum_{j=k}^{\infty}\beta_{j}\right)$ unoccupied space. At most $\alpha_{k}$ of the lines of order $k$ are more than half unoccupied by Lemma 14. Hence the maximum amount of unoccupied space in $H_{a+2}\cup\cdots\cup H_{L}$ is a solution to the following quadratic program
$\max\quad\quad$ $\sum_{k=1}^{\infty}\left(\alpha_{k}\beta_{k}+\frac{1}{2}\alpha_{k}\left(W-\sum_{j=k}^{\infty}\beta_{j}\right)\right)\quad\text{subject to},$
$0\leq\alpha_{k}\leq\frac{1}{2}h_{\text{OPT}}\quad\text{for all }k\in\mathbb{N},$
$\sum_{k=1}^{\infty}\alpha_{k}=h,$
$0\leq\beta_{k}\leq\frac{1}{2k+1}W\quad\text{for all }k\in\mathbb{N},$
$\sum_{k=1}^{\infty}\beta_{k}\leq W.$
@@
$\max\quad\quad\frac{1}{2}$ $\sum_{k=1}^{\infty}\alpha_{k}\gamma_{k}\quad\text{where }\gamma_{k}=\beta_{k}+W-\sum_{j=k+1}^{\infty}\beta_{j}.$
Let $k_{1}$ and $k_{2}$ be different indices such that $\gamma_{k_{1}}\geq\gamma_{k_{2}}\geq\gamma_{k}$ for all $k\in\mathbb{N}\setminus\{k_{1}\}$. Then setting $\alpha_{k_{1}}=\min\left\{\frac{1}{2}h_{\text{OPT}},h\right\}$, $\alpha_{k_{2}}=\max\left\{0,h-\frac{1}{2}h_{\text{OPT}}\right\}$ and $\alpha_{k}=0$ for $k\in\mathbb{N}\setminus\{k_{1},k_{2}\}$ maximizes the objective (provided that $\gamma_{k_{1}},\gamma_{k_{2}}\geq 0$), because by Lemma 11 we have that $h=y_{r_{L}}-y_{r_{a+1}}\leq h_{\text{max}}\leq h_{\text{OPT}}$, as $r_{B}$ is the leftmost bottom supporter of $r_{a+1}$. If $h\leq\frac{1}{2}h_{\text{OPT}}$, then $\alpha_{k_{2}}=0$ and the maximum objective value is $\frac{1}{2}h\gamma_{k_{1}}=\frac{2}{3}hW$ when $k_{1}=1$, $\beta_{1}=\frac{1}{3}W$ and $\beta_{k}=0$ for all $k\geq 2$. And observe that $\frac{2}{3}hW=\frac{1}{2}hW+\frac{1}{6}hW\leq\frac{1}{2}hW+\frac{1}{12}h_{\text{OPT}}W$ as $h\leq\frac{1}{2}h_{\text{OPT}}$. Otherwise, if $h>\frac{1}{2}h_{\text{OPT}}$, then the objective is $\frac{1}{4}h_{\text{OPT}}\gamma_{k_{1}}+\frac{1}{2}(h-\frac{1}{2}h_{\text{OPT}})\gamma_{k_{2}}$. There exists a maximum solution with $\beta_{k}=0$ for $k\not\in\{k_{1},k_{2}\}$, because $\beta_{k}>0$ only increases the objective when $\alpha_{k}>0$, which is only the case for $k\in\{k_{1},k_{2}\}$. Thus if $k_{1}<k_{2}$, then, as $h\leq h_{\text{OPT}}$, the objective becomes
$\frac{1}{4}h_{\text{OPT}}(\beta_{k_{1}}+W-\beta_{k_{2}})+\frac{1}{2}\left(h-\frac{1}{2}h_{\text{OPT}}\right)(\beta_{k_{2}}+W)$ $\leq\frac{1}{4}h_{\text{OPT}}\beta_{k_{1}}+\frac{1}{2}hW$
and in case $k_{1}>k_{2}$, the objective is
$\frac{1}{4}h_{\text{OPT}}(\beta_{k_{1}}+W)+\frac{1}{2}\left(h-\frac{1}{2}h_{\text{OPT}}\right)(\beta_{k_{2}}+W-\beta_{k_{1}})\leq\frac{1}{4}h_{\text{OPT}}\beta_{k_{2}}+\frac{1}{2}hW.$
This is maximum for $\min\{k_{1},k_{2}\}=1$ and $\beta_{\min\{k_{1},k_{2}\}}=\frac{1}{3}W$. This also gives the desired result $\frac{1}{12}h_{\text{OPT}}W+\frac{1}{2}hW$. ∎